# Trust-informed Decision-Making Through An Uncertainty-Aware Stacked Neural Networks Framework Case Study in COVID-19 Classification

### v1

@@ III-C Evaluation Metrics
Thus, new quantitative performance metrics have been proposed to evaluate trust-informed predictions (based on uncertainty estimations) by combining the outputs of both models, in a manner analogous to the traditional confusion matrix. In the context of binary classification, the base model’s predictions are compared with the ground truth labels, resulting in four categories: true positive (TP), false positive (FP), true negative (TN), and false negative (FN). Concurrently, the meta-model’s predicted confidence labels are compared with the ground truth trustworthy labels, yielding four additional categories: true trustworthy (TT), false trustworthy (FT), true untrustworthy (TU), and false untrustworthy (FU). By integrating these correctness and confidence classifications, 16 distinct outcomes are generated, as illustrated in Table I. For ease of reference, the proposed confusion matrix is called the trust-informed confusion matrix.
@@
Ground Truth Positive Trustworthy TPTT TPFU FNTT FNFU
@@
Considering the proposed confusion matrix, optimal performance is achieved when the model confidently makes correct predictions and is uncertain when an incorrect decision is likely. This means the model should never confidently make an incorrect prediction, and conversely, should never be uncertain when making a correct prediction. The idea is that when the model flags a prediction as uncertain, it should be reviewed by an expert for further examination. For the model to be efficient, only a small percentage of predictions—those where the model is unlikely to be accurate (i.e., where the model recognizes its own incorrectness)—should be flagged for further review. This approach minimizes the workload for expert review while maintaining high reliability.
Building upon the trust-informed confusion matrix, several quantitative criteria have been established as follows. These criteria are designed to evaluate how closely the model adheres to the optimal performance principles described above. These criteria are:
  • Confident Accuracy Rate (CAR): CAR is a metric that quantifies the proportion of instances where the model is accurate and confident, indicating its trustworthiness. It is defined as Eq. 8. The rationale behind naming this ratio Confident Accuracy lies in the definition of the meta-train set target variable as ”trustworthy,” as described in Eq. 1. In this context, ”trustworthy” signifies predictions that are both correct and made with high confidence. Generally, a higher CAR is desirable.
    $CAR=\frac{TPTT+TNTT}{Total\ Outcomes}$ (8)
  • Confident Precision Rate (CPR): CPR is a metric that measures the proportion of accurate predictions fagged as trustworthy, formulated by Eq. 9. Similarly, the rationale behind naming this ratio Confident Precision lies in the fact that this metric captures the percentage of correct predictions made with high confidence (which, according to the Eq. 1, are flagged as trustworthy) among all correct predictions. A CPR close to 1 is desirable, indicating that the model’s confident predictions are almost always correct.
    $CPR=\frac{TPTT+TNTT}{TAP}$ (9)
    where TAP stands as total accurate predictions defined by Eq. 10:
    $TAP=$ $\ TPTT+TPFU+TPFT+TPTU+$
    $\ TNTT+TNFU+TNFT+TNTU$ (10)
  • Trust Precision Rate (TPR): TPR is a metric that measures the proportion of predictions flagged as trustworthy that are actually trustworthy, formulated by Eq. 11. A TPR close to 1 is desirable. Naturally, $1-TPR$ represents the percentage of records that the model erroneously flags as trustworthy.
    $TPR=\frac{TPTT+TNTT}{TTP}$ (11)
@@
  • False Trust Rate (FTR): FTR measures the proportion of incorrect predictions that the model has erroneously classified as trustworthy. This measure is crucial in applications where false trust in incorrect predictions could lead to significant consequences. A lower FTR is desirable, as it indicates fewer instances where the model’s incorrect predictions are mistakenly trusted. The FTR is defined as Eq. 13:
    $FTR=\frac{FPFT+FNFT}{Total\ Outcomes}$ (13)
  • Review Alert Ratio (RAR): RAR is a metric that measures the proportion of predictions flagged as untrustworthy, requiring expert review, represented by Eq. 14. This metric encompasses all untrustworthy predictions, irrespective of whether they genuinely necessitate review or are incorrectly flagged as untrustworthy. A high RAR signifies inefficiency, indicating that a larger proportion of instances are being flagged for review, thereby increasing the review burden.
    $RAR=\frac{TUP}{Total\ Outcomes}$ (14)
@@
  • Miscalibration Review Ratio (MRR): MRR quantifies the proportion of predictions flagged for review due to the base model’s miscalibration. Specifically, it measures the rate at which the base model, despite being correct, has low confidence (indicating a miscalibration in the base model’s confidence assessment) and thus flags predictions as untrustworthy. A lower MRR indicates better calibration of the base model’s confidence with its actual predictive accuracy. The MRR is defined as Eq. 16:
    $MRR=\frac{TPTU+TNTU}{TUP}$ (16)
  • True Review Ratio (TRR): TRR is a metric that quantifies the proportion of instances where the model correctly identifies its incorrect predictions and flags them as untrustworthy formulated by Eq. 17. This metric indicates the model’s ability to recognize and appropriately flag its own errors. A TRR close to 1 is desirable, indicating that the model effectively flags only incorrect instances as untrustworthy.
    $TRR=\frac{FNTU+FPTU}{TUP}$ (17)
  • False Review Ratio (FRR): FRR quantifies the proportion of correct predictions that are incorrectly flagged as untrustworthy due to a meta-model error, thereby requiring unnecessary review, as formulated in Eq. 18. An FRR close to 0 is desirable.
    $FRR=\frac{TPFU+TNFU}{TUP}$ (18)
IV Data sets
@@ V-A Transfer Learning as Image Embedding Generator
The dataset utilized in this study, COVIDx CXR-4, lacks sufficient data to train a deep neural network effectively using CNNs from scratch. Therefore, this study adopted the transfer learning (TL) approach to prepare the dataset for the proposed method. TL is a robust technique within machine learning that alleviates the computational burden and extensive data requirements associated with training CNNs from the ground up.
Here, the primary goal of TL was to transfer knowledge by using pre-trained models as an embedding function (or, more simply, feature extractors). These models, initially trained and optimized on extensive datasets such as ImageNet, had their final fully-connected layers removed, and the remaining network layers were frozen. This repurposed the networks as fixed feature extractors for the COVIDx CXR-4 dataset, effectively harnessing their pre-established computational intelligence for new data applications.
@@ V-B Base model configuration
The embedding vectors (feature representations) extracted from the pre-trained models served as inputs for the base model. In this study, the base model is built using neural networks to perform binary classification. Considering the three different UQ techniques employed (MCD, Ensemble, and EMCD), the following configurations for the base model were devised and implemented:
V-B1 A Neural Network with MCD
Here, a fully connected layers with an output layer equipped with a softmax function is employed. To determine the optimal architecture, the dataset was first split into training and testing sets using a 70/30 ratio. The training set was used to train the models, while the test set was reserved for evaluating their performance. Various architectures were explored by employing Keras Tuner’s Hyperband algorithm. Keras Tuner, a library built on top of TensorFlow and Keras, automates the hyperparameter tuning process. Hyperband, an advanced tuning algorithm, dynamically allocates resources to train multiple models, stopping those that underperform early. The search space included variations in the number of hidden layers, ranging from one to four, and the number of neurons per layer, ranging from 16 to 512. Given the imbalanced nature of the dataset, class weights were computed to ensure balanced learning across classes. These weights were incorporated during the training process to prevent bias towards the majority class. The best-performing architecture was selected as the optimal base model for evaluating the proposed method. Additionally, MCD was employed during the prediction phase to estimate the model’s uncertainty, which was crucial for the subsequent stages of the study. In this study, the number of iterations for MCD was set to 100.
@@ V-B1 A Neural Network with MCD
V-B2 An Ensemble of Neural Networks
In this setting, the base model configuration leverages an ensemble approach. Thus, instead of relying on a single neural network, an ensemble of 30 distinct neural networks was employed. For each model in the ensemble, a neural network was built with a random number of hidden layers (ranging from one to four) and neurons per layer (ranging from 16 to 512) employing a softmax in the output layer. Similarly, class weights were utilized to address data imbalance for each neural network in the ensemble.
Table S2 in the Supplementary material provides a summary of the architectures for each model within the ensemble across the different pre-trained models.
V-B3 An Ensemble of Neural Networks with MCD
This setting integrates MCD with an ensemble of neural networks. The purpose of this configuration is to observe how the introduction of dropout during inference affects uncertainty estimation and model robustness. This approach ensures a direct comparison with the previous ensemble setup by using the same ensemble model configuration as the prior setting, with the addition of MCD during the inference phase.
@@ V-C Meta-model configuration
The meta-model is designed as a neural network, and its optimal architecture is determined using the Keras Tuner’s Hyperband algorithm. This method, similar to that employed for the single neural network with the MCD base model, efficiently explores the hyperparameter space to identify the most effective architecture. The search space for the meta-model includes variations in the number of hidden layers, ranging from one to four, and the number of neurons per layer, ranging from 16 to 512, allowing the model to adapt to the complexities of the data. Class weights are again utilized to address data imbalance, maintaining fairness in the learning process.
For the meta-model dataset generation, the confidence threshold is a crucial parameter that impacts the labeling of the data. In this study, five different thresholds were used: 0.05, 0.1, 0.2, 0.3, and 0.4. These thresholds determine the cut-off points for labeling predictions as confident or not, which in turn affects the training and performance of the meta-model. By experimenting with these different thresholds, the study aims to identify the optimal level at which the model can reliably distinguish between confident and uncertain predictions.
@@ VI Results and Discussion
This section delves into the empirical outcomes derived from the application of the proposed method:
  • Initially, the analysis commences with a focus on a selected confidence threshold of 0.1, evaluating the performance across three distinct UQ settings: MCD, Ensemble, and EMCD. This section discusses their respective contributions to enhancing trust-informed decision-making. Additionally, this examination is broadened to include a discussion about the effects of various pre-trained models on the models’ performance, providing a deeper understanding of how foundational architectures influence model reliability.
  • Subsequently, the discourse broadens to include a comparative analysis across a spectrum of confidence thresholds, ranging from 0.05 to 0.4. This section aims to delineate how varying confidence levels influence the model’s accuracy and the efficacy of trust-informed decision-making, providing more insights into the model’s robustness.
  • The final section investigates the integration of PE alongside image embeddings as inputs to the meta-models. This exploration is aimed at discerning the impact of such integrations on the overall efficacy and accuracy of the meta-model, thereby shedding light on the potential enhancements introduced by combining these inputs.
VI-A Results and discussion at Confidence Threshold 0.1:
The performance evaluation process begins with partitioning the dataset into training and testing sets through stratified sampling based on the target variable. This ensures that both sets maintain a similar distribution of classes. While evaluating a model on a single test set might provide an initial performance snapshot, it does not guarantee reproducible results across different data splits. This is because each random split could present the model with different distributions of data. When a model is trained on a specific training set, it learns patterns based on the distribution of that set. Similarly, the testing set used for evaluation will also have its unique distribution. Therefore, repeating the process with different splits can yield varying results, highlighting the non-reproducibility of single evaluations.
@@
To address this variability and better gauge the model’s generalization capabilities, the training and evaluation process is repeated 30 times. For each iteration, the dataset is randomly split into training and testing sets, with a test size randomly chosen between [20%- 40%] of the data. This repetitive approach ensures that the model is exposed to various data distributions, enable evaluation of the model’s ability to generalize across various potential data distributions.
The evaluation metrics in this study are twofold. For the base model and meta-model, metrics include the F1 score and the Area Under the Curve (AUC). The F1 score and AUC are particularly suited for this study due to the imbalanced nature of the dataset.
Table II summarizes the base model performance across three UQ settings as well as pre-trained models. The evaluation of base model performance across different UQ settings reveals distinct patterns in classification effectiveness. The Ensemble setting, which utilizes 30 ANNs with varying architectures, generally improves F1 scores compared to MCD, reflecting enhanced precision and recall due to the collective decision-making of multiple networks. However, this setting shows a tendency for lower AUC values, likely due to the inclusion of networks that vary in strength, potentially affecting overall class discrimination. The EMCD setting, which applies dropout during inference across the ensemble, exhibits performance metrics similar to the Ensemble setting. This similarity is logical, as EMCD combines the architectural diversity of the ensemble with dropout at inference, maintaining the ensemble’s overall classification characteristics. A higher AUC for MCD suggests that fine-tuning the single neural network contributed to better class discrimination, as evidenced by the AUC. In contrast, the Ensemble and EMCD settings improve precision and recall (reflected in higher F1 scores) but introduce variability due to the random architectures, which may affect the overall AUC. Results across pre-trained models highlight ViT’s strength in leveraging its architectural design to enhance classification accuracy and discrimination. The BiT model also shows competitive performance, particularly in F1 scores, reflecting its capacity to balance precision and recall effectively. However, BiT’s AUC values are slightly lower than those of ViT, indicating some limitations in its ability to separate classes with the same level of distinction. EfficientNetB0 exhibits the lowest performance among the pre-trained models, with both F1 scores and AUC values trailing behind ViT and BiT. This suggests that while EfficientNetB0 can achieve reasonable classification results, it may not capture the nuances of the data as effectively as the other models. This performance gap behind ViT and BiT could be attributed to the architecture of EfficientNetB0, which is a lighter model with significantly fewer parameters compared to ViT and BiT. These findings suggest that selecting ViT with MCD offers the finest balance of accuracy and robustness, making it the preferred choice for the problem addressed in this study.
@@
ViT Ensemble %88.0198 ± 0.2486 %84.8285 ± 0.2681
ViT EMCD %87.0202 ± 0.4014 %84.1501 ± 0.5857
@@
ViT Ensemble %97.6332 ± 0.2767 %99.7606 ± 0.0379
ViT EMCD %97.4426 ± 0.3090 %99.6196 ± 0.0865
Similarly, Table III illustrates the meta-model performance across three UQ settings as well as pre-trained models. It is worth noting that the meta-model across all pre-trained models and UQ settings is a single neural network fine-tuned using the Keras Tuner. The influence of the UQ settings lies in the additional input variable, Prediction Entropy (PE), which is derived from the uncertainty estimation of each UQ method, as well as the target variable defined by Eq. 1, which also depends on PE. Overall, all models exhibit high AUC values, consistently surpassing 99%, indicating a strong ability to discriminate between classes across the UQ settings and pre-trained models. Notably, the ViT models achieved the highest F1 scores and AUC values. In comparison, the EfficientNetB0 models show slightly lower F1 scores but still maintain high AUC values, reflecting reliable classification performance with minor variations across different UQ methods. The BiT models perform consistently across all UQ settings, although they exhibit slightly lower F1 scores and AUC values compared to the ViT models, suggesting that ViT models may be better suited for the task at this confidence threshold. Among the UQ methods, the Ensemble method provides the best performance in terms of AUC, particularly with the ViT models. The MCD method also performs competitively, especially with the ViT models, though it falls slightly behind the Ensemble method in terms of F1 scores. Meanwhile, the EMCD method, while maintaining high AUC values, shows slightly lower F1 scores, indicating that it may not balance precision and recall as effectively as the other methods at this specific threshold.
((a)) CAR
((b)) CPR
((c)) TPR
((d)) FTR
((e)) RAR
((f)) MRR
((g)) TRR
((h)) FRR

### v2

@@ III-C2 Metric Formulation for the U-SNN Approach
The proposed U-SNN method employs a dedicated meta-model trained specifically to discern whether a prediction from the base model is trustworthy or requires additional review. In a manner analogous to the traditional confusion matrix, the base model’s predictions are compared with the ground truth labels, resulting in four categories: true positive (TP), false positive (FP), true negative (TN), and false negative (FN). Concurrently, the meta-model’s predicted confidence labels are compared with the ground truth trustworthy labels, yielding four additional categories: true trustworthy (TT), false trustworthy (FT), true untrustworthy (TU), and false untrustworthy (FU). By integrating these correctness and confidence classifications, 16 distinct outcomes are generated, as illustrated in Table I. For ease of reference, the proposed confusion matrix is called the trust-informed confusion matrix.
@@
Ground
Truth
Positive Trustworthy TPTT TPFU FNTT FNFU
@@
From this expanded structure, the evaluation metrics are reformulated as follows:
  • CR: Calculated by measuring the proportion of predictions classified as trustworthy, irrespective of correctness, denoted by Eq. 9. Importantly, all trustworthy predictions—regardless of their correctness—are conceptually equivalent to confident predictions in the threshold-based method.
    $CR=\frac{TTP}{Total\ Outcomes}$ (9)
@@
  • FCR: Measured by the proportion of false trustworthy predictions relative to all predictions, presented by Eq. 11. Notably, false trustworthy predictions are conceptually equivalent to false confident outcomes in the context of the threshold-based method.
    $FCR=\frac{FPFT+FNFT}{Total\ Outcomes}$ (11)
  • CE: Quantified by the proportion of false trustworthy predictions among all trusted predictions, as calculated by Eq. 12. Importantly, the proportion of false trustworthy predictions among all trusted predictions conceptually corresponds to the proportion of false certain predictions among all confident predictions in the context of the threshold-based method.
    $CE=\frac{FPFT+FNFT}{TTP}$ (12)
  • UR: Calculated by the proportion of predictions classified as untrustworthy (flagged for review) relative to all predictions, as defined in Eq. 13.
    $UR=\frac{TUP}{Total\ Outcomes}$ (13)
@@
  • RR: Quantifies the proportion of flagged predictions for review that were unnecessary, as calculated by Eq 15. In the U-SNN method, redundant referrals arise when predictions correctly classified by the base model are mistakenly flagged as untrustworthy by the meta-model. These include correctly predicted outcomes incorrectly labeled as untrustworthy (TPFU, TNFU) and correctly predicted outcomes flagged as uncertain (TPTU, TNTU). Importantly, TPTU and TNTU are conceptually equivalent to false uncertainty (FU) in the context of the threshold-based method, where correct predictions are unnecessarily flagged for review.
    $RR=\frac{TPFU+TNFU+TPTU+TNTU}{TUP}$ (15)
IV Data set
@@ V-A Transfer Learning as Image Embedding Generator
This study adopted the transfer learning (TL) approach to prepare the dataset for the proposed method. Here, the primary goal of TL was to transfer knowledge by using pre-trained models as an embedding function (or, more simply, feature extractors). These models, initially trained and optimized on extensive datasets such as ImageNet, had their final fully-connected layers removed, and the remaining network layers were frozen. This repurposed the networks as fixed feature extractors for the COVIDx CXR-4 dataset, effectively harnessing their pre-established computational intelligence for new data applications.
@@ V-B Base model configuration
The embedding vectors (feature representations) extracted from the pre-trained models served as inputs for the base model. In this study, the base model is built using neural networks to perform binary classification. Considering the UQ technique employed in this study, a fully connected layers with an output layer equipped with a softmax function is employed. To determine the optimal architecture, the train set was first split into training and evaluation sets using a 70/30 ratio. The training set was used to train the models, while the evaluation set was reserved for evaluating their performance. Various architectures were explored by employing Keras Tuner’s Hyperband algorithm. Keras Tuner, a library built on top of TensorFlow and Keras, automates the hyperparameter tuning process. Hyperband, an advanced tuning algorithm, dynamically allocates resources to train multiple models, stopping those that underperform early. The search space included variations in the number of hidden layers, ranging from one to four, and the number of neurons per layer, ranging from 16 to 512. Given the imbalanced nature of the dataset, class weights were computed to ensure balanced learning across classes. These weights were incorporated during the training process to prevent bias towards the majority class. The best-performing architecture was selected as the optimal base model for evaluating the proposed method. Additionally, MCD was employed during the prediction phase to estimate the model’s uncertainty. In this study, the number of iterations for MCD was set to 100.
@@
For a fair and unbiased comparison, both the traditional threshold-based method and the proposed U-SNN utilized the same trained base model.
@@ V-C Meta-model configuration
The meta-model is designed as a neural network, and its optimal architecture is determined using the Keras Tuner’s Hyperband algorithm. This method, similar to that employed for the base model, efficiently explores the hyperparameter space to identify the most effective architecture. The search space for the meta-model includes variations in the number of hidden layers, ranging from one to four, and the number of neurons per layer, ranging from 16 to 512, allowing the model to adapt to the complexities of the data. Class weights are again utilized to address data imbalance, maintaining fairness in the learning process.
For the meta-model dataset generation, the confidence threshold is a crucial parameter that impacts the labeling of the data. In this study, five different thresholds were used: 0.05, 0.1, 0.2, and 0.4. These thresholds determine the cut-off points for labeling predictions as trustworthy or not, which in turn affects the training and performance of the meta-model.
@@ VI Results and Discussion
Evaluating a model on a single test set might provide an initial performance snapshot, it does not guarantee reproducible results across different data splits. To address this variability and better gauge the model’s generalization capabilities, the training and evaluation process is repeated 30 times. For each iteration, the dataset is randomly split into training and testing sets, with a test size randomly chosen between [20%- 40%] of the data. This repetitive approach ensures that the model is exposed to various data distributions, enable evaluation of the model’s ability to generalize across various potential data distributions.
The classification performance of the base models, assessed independently of the uncertainty evaluation framework, is summarized in Table II. Three different pre-trained architectures—BiT, EfficientNetB0, and ViT—were evaluated based on their respective F1 scores and the AUC. Among the tested models, ViT yielded the highest classification performance, achieving an average F1 score of 86.95% (±0.46) and an AUC of 94.21% (±0.32). The BiT model followed closely, demonstrating an F1 score of 85.75% (±0.40) and an AUC of 92.60% (±0.33). EfficientNetB0 exhibited slightly lower performance, with an F1 score of 84.45% (±0.46) and an AUC of 91.58% (±0.23). These results confirmed that all base models provided strong predictive capabilities, suitable as foundational classifiers for subsequent uncertainty analysis. Notably, the consistently high AUC values across all three models indicate robust discrimination ability, further supporting their effectiveness in the underlying classification task.
@@
As the proposed U-SNN approach incorporates an additional meta-model specifically designed to assess the trustworthiness of predictions, the overall reliability of the U-SNN directly depends on the performance of this meta-model. Accordingly, before comparing the uncertainty performance metrics between the U-SNN and the traditional threshold-based method, the classification performance of the meta-model was evaluated separately. This evaluation aimed to verify the meta-model’s capability to correctly distinguish trustworthy predictions (correctly identifying correct and confident predictions) from untrustworthy predictions (either incorrect or uncertain). It should be noted that these results pertain exclusively to the proposed U-SNN, as the traditional threshold-based method does not include a meta-model. The performance of the meta-model is influenced by the confidence threshold, as determined by the target variable definition in Equation 1. Table III illustrates the meta-model performance across three as pre-trained models at confidence threshold 0.1. Supplementary Table S.2 presents a summary of the meta-model’s error-based performance results across confidence thresholds of 0.05, 0.2, and 0.4.
Across the three pre-trained architectures examined—BiT, EfficientNetB0, and ViT—the meta-model consistently demonstrated robust performance, with high F1 scores and exceptionally high AUC values. Specifically, the ViT-based meta-model yielded the highest performance, achieving an F1 score of 97.48% (±0.32) and an AUC of 99.63% (±0.11). The BiT-based meta-model closely followed, recording an F1 score of 96.65% (±0.35) and an AUC of 99.50% (±0.09), while the EfficientNetB0-based meta-model displayed slightly lower but still strong results, with an F1 score of 96.37% (±0.45) and an AUC of 99.47% (±0.13). These results indicate that the meta-model reliably distinguishes between trustworthy and untrustworthy predictions.
Following, a direct comparison of uncertainty metrics between the U-SNN framework and the traditional threshold-based approach is presented in Table IV.
@@
To start, let’s first examine the results at confidence threshold 0.1. Figure 5 illustrates a comparative analysis of the U-SNN and the threshold-based method at a confidence threshold of 0.1. At a confidence threshold of 0.1, both the U-SNN and threshold-based approaches displayed similar CR across all models. This observation suggests that both methods were equally assertive in labeling predictions as confident. However, similarity in CR values does not inherently indicate reliability, as it does not account for the correctness of these confident predictions. The distinction between the two methods becomes evident when examining the FCR and CE. FCR measures how often is the model confidently incorrect among all predictions. The U-SNN consistently demonstrated a substantial reduction in FCR compared to the threshold-based approach. For the BiT model, U-SNN achieving a 55.4% reduction in FCR and marking a 48.4% and 47.7% improvement in FCR for EfficientNetB0 and ViT, respectively. This indicates that U-SNN was considerably more effective in preventing confidently incorrect predictions. The implication is significant: in high-stakes scenarios, such as medical diagnosis, confidently incorrect predictions can mislead operators into trusting erroneous results. By substantially reducing FCR, the U-SNN framework mitigates this critical risk.
Similarly, the CE metric further highlights the robustness of U-SNN by assessing the proportion of errors among confident predictions. U-SNN achieved a 55.9% reduction in CE for the BiT model, along with reductions of 50.9% and 48.0% for EfficientNetB0 and ViT, respectively. These improvements indicate that when U-SNN flags a prediction as confident, there is a significantly higher likelihood that it is indeed correct. In contrast, the higher CE in the threshold-based approach reflects a vulnerability, where confidently incorrect predictions are more prevalent, potentially eroding trust in the system.
The UR remained relatively similar between the methods, suggesting that both approaches referred a similar proportion of predictions for further review. However, the key distinction arises in the RR metric, which captures the inefficiency of unnecessary reviews: For BiT, RR dropped from 97.40% to 85.92%, a 11.8% reduction in unnecessary referrals. For EfficientNetB0 reflecting a 12.3% improvement and For ViT representing a 13.9% reduction in RR. Although UR rates were similar, the substantial reduction in RR demonstrates that U-SNN was far more selective and efficient in its referrals. This has significant practical implications: unnecessary referrals increase workload, delay decision-making, and strain resources. By reducing RR, the U-SNN approach enhances operational efficiency, ensuring that human reviewers focus only on genuinely uncertain or complex cases that require further analysis.
TABLE IV: Comparison of traditional threshold-base model and proposed method (U-SNN)
Pretrained Confidence Threshold Approach CR FCR CE UR $R$
BiT 0.05 U-SNN 29.4514 ± 4.4865 0.2081 ± 0.0749 0.6902 ± 0.1652 70.5486 ± 4.4865 87.4771 ± 0.6375
Threshold-Based 29.4052 ± 4.6884 0.7444 ± 0.2518 2.5239 ± 0.7174 70.5948 ± 4.6884 97.4181 ± 0.6319
EfficientNetB0 U-SNN 27.4149 ± 3.7112 0.2182 ± 0.0897 0.7724 ± 0.2322 72.5851 ± 3.7112 86.7379 ± 0.5236
Threshold-Based 26.9223 ± 4.3993 0.7121 ± 0.2046 2.6735 ± 0.7823 73.0777 ± 4.3993 97.2557 ± 0.5887
ViT U-SNN 39.3139 ± 5.8126 0.2212 ± 0.0981 0.5408 ± 0.1706 60.6861 ± 5.8126 86.4584 ± 1.0562
Threshold-Based 39.8193 ± 5.1586 0.7203 ± 0.2274 1.7866 ± 0.4357 60.1807 ± 5.1586 98.1765 ± 0.4355
BiT 0.1 U-SNN 38.7992 ± 5.0042 0.4395 ± 0.1458 1.107 ± 0.2486 61.2008 ± 5.0042 85.9213 ± 0.828
Threshold-Based 39.1187 ± 4.9228 0.9857 ± 0.2939 2.5228 ± 0.6806 60.8813 ± 4.9228 97.4005 ± 0.6731
EfficientNetB0 U-SNN 36.1502 ± 3.9059 0.4755 ± 0.1488 1.2907 ± 0.2857 63.8498 ± 3.9059 85.3193 ± 0.6299
Threshold-Based 35.7458 ± 4.5747 0.9323 ± 0.2528 2.6277 ± 0.7215 64.2542 ± 4.5747 97.2246 ± 0.5881
ViT U-SNN 48.3345 ± 5.5188 0.4623 ± 0.1625 0.9315 ± 0.2372 51.6655 ± 5.5188 84.5481 ± 1.2343
Threshold-Based 48.841 ± 4.9201 0.8828 ± 0.2492 1.7931 ± 0.413 51.159 ± 4.9201 98.1729 ± 0.459
BiT 0.2 U-SNN 52.9509 ± 5.2666 1.1213 ± 0.2945 2.0862 ± 0.3533 47.0491 ± 5.2666 83.098 ± 1.1284
Threshold-Based 53.8595 ± 4.8049 1.3808 ± 0.3561 2.5623 ± 0.608 46.1405 ± 4.8049 97.4307 ± 0.7138
EfficientNetB0 U-SNN 50.0759 ± 4.1463 1.2038 ± 0.2583 2.3804 ± 0.339 49.9241 ± 4.1463 82.6654 ± 0.8199
Threshold-Based 50.18 ± 4.3741 1.3476 ± 0.3392 2.7034 ± 0.7264 49.82 ± 4.3741 97.2567 ± 0.5432
ViT U-SNN 60.0983 ± 4.7476 1.0308 ± 0.2663 1.692 ± 0.3233 39.9017 ± 4.7476 81.4207 ± 1.411
Threshold-Based 60.679 ± 4.5586 1.1108 ± 0.2762 1.8227 ± 0.3947 39.321 ± 4.5586 98.205 ± 0.5112
BiT 0.4 U-SNN 75.6858 ± 3.5126 3.3056 ± 0.4536 4.3513 ± 0.4165 24.3142 ± 3.5126 76.3087 ± 1.4808
Threshold-Based 76.7593 ± 3.5632 1.9887 ± 0.4471 2.5922 ± 0.5726 23.2407 ± 3.5632 97.5396 ± 0.7978
EfficientNetB0 U-SNN 72.9876 ± 3.8683 3.4535 ± 0.4571 4.7129 ± 0.3947 27.0124 ± 3.8683 76.2099 ± 1.5066
Threshold-Based 74.1529 ± 3.4815 1.9709 ± 0.4456 2.6669 ± 0.6355 25.8471 ± 3.4815 97.1108 ± 0.5627
ViT U-SNN 77.2759 ± 3.3395 2.669 ± 0.4377 3.4373 ± 0.4289 22.7241 ± 3.3395 74.6181 ± 1.7885
Threshold-Based 78.0048 ± 3.1074 1.4212 ± 0.3269 1.8194 ± 0.3991 21.9952 ± 3.1074 98.2178 ± 0.5568
((a)) CR
((b)) FCR
((c)) CE
((d)) UR
((e)) $R$